OpenStudy (anonymous):

Solve: 24/2x-1=72/2x+1

OpenStudy (anonymous):

x=-1/12 :)

OpenStudy (anonymous):

24(2x+1)=72(2x-1) ........48x+24=144x-72 .........48x-144x=-72-24 ........x=1